The ratio of the cost of a DVD player to the cost of a movie is 9:2.
A DVD player costs $63 more than a movie.
What is the cost of a DVD player?

Answer by ikleyn(52818)

.
9x - 2x = 63


7x = 63


x = 63/7 = 9.


ANSWER.  2*9 = 18 dollars and 9*9 = 81 dollars.